# A unified relative entropy framework for macroscopic limits of Vlasov–Fokker–Planck equations

### v1

@@ 3.1. Bounded Lipschitz distance: connection with entropy and dissipation
that is, $\bar{\rho}$ is continuous in the duality with continuous bounded functions, for a Borel vector field $\bar{u}$ satisfying
@@ 3.2. Modulated energy for the Riesz interaction
A central ingredient in the relative entropy framework is the control of the nonlocal interaction term. To this end, we employ modulated interaction energies, which provide a quantitative way of comparing two densities evolving under continuity equations. Such estimates were developed in [24, Lemma 2.1] (see also [27, 50]), and we recall the version adapted to our setting below.
@@ 3.3. Relative entropy estimate
$\quad=-\frac{\rm B}{{\rm A}}\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}f\left((u-\xi)\otimes(u-\xi)-I_{d}\right):\nabla u\,dxd\xi-\frac{1}{{\rm A}\tau}\int_{\mathbb{R}^{d}}\rho_{f}u_{f}\cdot(u_{f}-u)\,dx$
@@
$\quad-\frac{1}{{\rm A}\tau}\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\frac{1}{f}|\nabla_{\xi}f+(\xi-u_{f})f|^{2}dxd\xi+\frac{1}{{\rm A}\tau}\int_{\mathbb{R}^{d}}\rho_{f}u_{f}\cdot(u_{f}-u)\,dx$
@@
$II_{1}=\frac{\rm B}{\rm A}\int_{\mathbb{R}^{d}}\frac{\rho_{f}}{\rho}\nabla\cdot(\rho u)\,dx=\frac{\rm B}{\rm A}\int_{\mathbb{R}^{d}}\rho_{f}u\cdot\nabla\log\rho\,dx+\frac{\rm B}{\rm A}\int_{\mathbb{R}^{d}}\rho_{f}I_{d}:\nabla u\,dx$
@@
$=-\frac{\rm B}{{\rm A}}\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}u\otimes(u-\xi)f:\nabla u\,dxd\xi-\frac{\rm B}{{\rm A}}\int_{\mathbb{R}^{d}}\rho_{f}(u-u_{f})\cdot\nabla\log\rho\,dx$
@@
$II$ $=\frac{\rm B}{\rm A}\int_{\mathbb{R}^{d}}\rho_{f}u_{f}\cdot\nabla\log\rho\,dx+\frac{\rm B}{\rm A}\int_{\mathbb{R}^{d}}\rho_{f}I_{d}:\nabla u\,dx-\frac{\rm B}{{\rm A}}\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}u\otimes(u-\xi)f:\nabla u\,dxd\xi$
@@
$-\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\left(f(u-\xi)\otimes(u-\xi)-\rho_{f}I_{d}\right):\nabla u\,dxd\xi$
@@ 4. Quantitative limits: Diffusive, High-Field, and gSQG
Building on the quantitative stability tools developed in Section 3, we now pass to three hydrodynamic regimes of the kinetic model: the diffusive, high-field, and gSQG limits. Each regime corresponds to a different balance between transport, interaction, and relaxation, yet all can be analyzed within the same entropy-dissipation framework. The common strategy is to reformulate the limit equation in conservative form so that it fits into the structure of Proposition 3.2, thereby enabling a comparison with the kinetic dynamics via relative entropy. In this way, the relative entropy is coupled with BL-type controls or modulated interaction energies for the nonlocal term, and the Fisher-information dissipation can be fully exploited to handle mildly well-prepared data.
A unifying feature across these limits is that the dissipation acts strongly enough to suppress microscopic fluctuations and transfer control to macroscopic quantities. This makes it possible to derive quantitative convergence in weak topologies while still obtaining strong stability of hydrodynamic observables in the well-prepared setting. The BL distance provides a bridge between entropy and weak convergence of measures, ensuring stability of densities under limited information, whereas modulated interaction energies tailored to the Riesz potential yield a comparison principle for evolving densities that accounts for the nonlocal structure. Together, these ingredients translate microscopic dissipation into macroscopic stability estimates that are effective even when initial data are only partially compatible with the limiting dynamics.
@@ 4.1. Diffusive limit
Formally, as stated in Section 1.2.1, one obtains the nonlinear diffusion equation as $\varepsilon\to 0$
@@
$\left|\int_{\mathbb{R}^{d}}\rho^{\varepsilon}({\rm u_{\varepsilon}}-u^{\varepsilon})\cdot{\rm e}_{\varepsilon}\,dx\right|\leq\frac{1}{2\varepsilon^{2}}\int_{\mathbb{R}^{d}}\rho^{\varepsilon}|u^{\varepsilon}-{\rm u_{\varepsilon}}|^{2}\,dx+\frac{\varepsilon^{2}}{2}\|{\rm e}_{\varepsilon}\|_{L^{\infty}}\|\rho^{\varepsilon}\|_{L^{1}}\leq\frac{1}{2\varepsilon^{2}}\int_{\mathbb{R}^{d}}\rho^{\varepsilon}|u^{\varepsilon}-{\rm u_{\varepsilon}}|^{2}\,dx+C\varepsilon^{3}$
@@
$\mathscr{H}[f^{\varepsilon}|M_{\rho,{\rm u_{\varepsilon}}}]\leq 2\mathscr{H}[f^{\varepsilon}|M_{\rho,0}]+C\varepsilon^{2}$
@@
$\mathscr{H}[f^{\varepsilon}|M_{\rho,0}]\leq 2\mathscr{H}[f^{\varepsilon}|M_{\rho,{\rm u_{\varepsilon}}}]+C\varepsilon^{2}$
@@
Case of well-prepared initial data. Estimate (4.6), together with Lemma 2.2, ensures the convergence of $f^{\varepsilon}$ and $\rho^{\varepsilon}$. For the momentum, we note
@@
Case of mildly well-prepared initial data. In contrast to the well-prepared case, here the initial data are only assumed to be close to equilibrium in a weak sense. To capture convergence in this regime, we work with the BL distance, which is well-suited to weak topologies and interacts effectively with the dissipation structure identified earlier.
From Lemmas 3.1 and 3.2, we obtain
$\sup_{0\leq t\leq T}{\rm d}^{2}_{\rm BL}(\rho^{\varepsilon}(t),\rho(t))+\int_{0}^{T}{\rm d}^{2}_{\rm BL}(f^{\varepsilon},M_{\rho,{\rm u_{\varepsilon}}})\,dt$
$\quad\leq C{\rm d}^{2}_{\rm BL}(\rho^{\varepsilon}_{0},\rho_{0})+C\int_{0}^{T}\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\frac{1}{f^{\varepsilon}}|\nabla_{\xi}f^{\varepsilon}+(\xi-u^{\varepsilon})f^{\varepsilon}|^{2}\,dxd\xi dt+C\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho^{\varepsilon}|u^{\varepsilon}-{\rm u_{\varepsilon}}|^{2}\,dxdt.$
The last two terms on the right-hand side are precisely the coercive contributions already controlled by the relative entropy method. Indeed, estimate (4.6) yields
$\int_{0}^{T}\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\frac{1}{f^{\varepsilon}}|\nabla_{\xi}f^{\varepsilon}+(\xi-u^{\varepsilon})f^{\varepsilon}|^{2}\,dxd\xi dt+\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho^{\varepsilon}|u^{\varepsilon}-{\rm u_{\varepsilon}}|^{2}\,dxdt$
$\quad\leq C\varepsilon^{2}\left(\mathscr{H}[f^{\varepsilon}|M_{\rho,0}](0)+\mathscr{P}[\rho^{\varepsilon}|\rho](0)\right)+C\varepsilon^{4}.$
Thus, we have
$\sup_{0\leq t\leq T}{\rm d}^{2}_{\rm BL}(\rho^{\varepsilon}(t),\rho(t))+\int_{0}^{T}{\rm d}^{2}_{\rm BL}(f^{\varepsilon},M_{\rho,{\rm u_{\varepsilon}}})\,dt$
$\quad\leq C{\rm d}^{2}_{\rm BL}(\rho^{\varepsilon}_{0},\rho_{0})+C\varepsilon^{2}\left(\mathscr{H}[f^{\varepsilon}|M_{\rho,0}](0)+\mathscr{P}[\rho^{\varepsilon}|\rho](0)\right)+C\varepsilon^{4}.$
At this point, it is convenient to compare the kinetic solution with the simpler Maxwellian $M_{\rho,0}$ rather than $M_{\rho,{\rm u_{\varepsilon}}}$. For any $\varphi\in W^{1,\infty}(\mathbb{R}^{d}\times\mathbb{R}^{d})$ we decompose

### v2

@@ 3.1. Weak-topology estimates: BL distance and negative Sobolev norms
that is, for every $\varphi\in C_{b}(\mathbb{R}^{d})$ the map
$t\mapsto\int_{\mathbb{R}^{d}}\varphi(x)\bar{\rho}(t,x)\,dx$
is continuous. Assume that the Borel vector field $\bar{u}$ satisfies
@@
The BL estimate above can be complemented by a negative Sobolev estimate, obtained by testing the same decomposition against $H^{\sigma}$ functions rather than bounded Lipschitz test functions.
Corollary 3.1.
Under the assumptions of Lemma 3.2, for each $\sigma>d$, there exists $C>0$ depending only on $d$ and $T$ such that for all $t\in[0,T]$,
$\int_{0}^{t}\|f-M_{\rho,u}\|_{H_{x,\xi}^{-\sigma}}^{2}\,ds$ $\leq C\int_{0}^{t}\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\frac{1}{f}|\nabla_{\xi}f+(\xi-u_{f})f|^{2}\,dxd\xi ds+C\int_{0}^{t}\int_{\mathbb{R}^{d}}\rho_{f}|u_{f}-u|^{2}\,dxds$
$\quad+C\|u\|_{L^{\infty}((0,T)\times\mathbb{R}^{d})}^{2}\int_{0}^{t}\|\rho_{f}-\rho\|_{L^{1}}^{2}\,ds+C\int_{0}^{t}\|\rho_{f}-\rho\|_{H^{-\sigma}}^{2}\,ds.$
In particular, if $u=0$, then
$\int_{0}^{t}\|f-M_{\rho,0}\|_{H_{x,\xi}^{-\sigma}}^{2}\,ds\leq C\int_{0}^{t}\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\frac{1}{f}|\nabla_{\xi}f+\xi f|^{2}\,dxd\xi ds+C\int_{0}^{t}\|\rho_{f}-\rho\|_{H^{-\sigma}}^{2}\,ds.$
Proof.
We follow the decomposition used in the proof of Lemma 3.2, namely,
$\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\varphi(x,\xi)\bigl(f-M_{\rho,u}\bigr)\,dxd\xi=I+II+III,$
where
$I:=\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\varphi(f-M_{\rho_{f},u_{f}})\,dxd\xi,\quad II:=\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\varphi(M_{\rho_{f},u_{f}}-M_{\rho_{f},u})\,dxd\xi,$
and
$III:=\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\varphi(M_{\rho_{f},u}-M_{\rho,u})\,dxd\xi.$
Fix $\varphi\in H^{\sigma}(\mathbb{R}^{d}\times\mathbb{R}^{d})$ with $\sigma>d$. Since $H^{\sigma}(\mathbb{R}^{d}\times\mathbb{R}^{d})\hookrightarrow L^{\infty}(\mathbb{R}^{d}\times\mathbb{R}^{d})$, the estimates for $I$ and $II$ obtained in Lemma 3.2 remain valid with $\|\varphi\|_{L^{\infty}}$ replaced by $C\|\varphi\|_{H^{\sigma}}$. Thus,
$|I|\leq C\|\varphi\|_{H^{\sigma}}\left(\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\frac{1}{f}|\nabla_{\xi}f+(\xi-u_{f})f|^{2}\,dxd\xi\right)^{1/2}$
and
$|II|\leq C\|\varphi\|_{H^{\sigma}}\left(\int_{\mathbb{R}^{d}}\rho_{f}|u_{f}-u|^{2}\,dx\right)^{1/2}.$
For $III$, we write
$III$ $=\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\varphi(x,\xi)(\rho_{f}-\rho)(x)(M_{1,u(x)}(\xi)-M_{1,0}(\xi))\,dxd\xi+\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\varphi(x,\xi)(\rho_{f}-\rho)(x)M_{1,0}(\xi)\,dxd\xi.$
Using the mean-value theorem in $u$, we have
$|M_{1,u}-M_{1,0}|\leq C|u|(1+|\xi|)M_{1,\theta u}$
for some $\theta\in(0,1)$, and hence
$\left|\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\varphi(\rho_{f}-\rho)(M_{1,u}-M_{1,0})\,dxd\xi\right|\leq C\|\varphi\|_{H^{\sigma}}\|u\|_{L^{\infty}}\|\rho_{f}-\rho\|_{L^{1}}.$
For the second term, by duality between $H_{x}^{-\sigma}$ and $H_{x}^{\sigma}$,
$\left|\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\varphi(x,\xi)(\rho_{f}-\rho)(x)M_{1,0}(\xi)\,dxd\xi\right|\leq\|\rho_{f}-\rho\|_{H_{x}^{-\sigma}}\left\|\int_{\mathbb{R}^{d}}\varphi(\cdot,\xi)M_{1,0}(\xi)\,d\xi\right\|_{H_{x}^{\sigma}}.$
Moreover, by Minkowski’s inequality and Cauchy–Schwarz inequality,
$\left\|\int_{\mathbb{R}^{d}}\varphi(\cdot,\xi)M_{1,0}(\xi)\,d\xi\right\|_{H_{x}^{\sigma}}$ $=\left\|\int_{\mathbb{R}^{d}}(1-\Delta_{x})^{\sigma/2}\varphi(\cdot,\xi)M_{1,0}(\xi)\,d\xi\right\|_{L_{x}^{2}}$
$\leq\int_{\mathbb{R}^{d}}\|(1-\Delta_{x})^{\sigma/2}\varphi(\cdot,\xi)\|_{L_{x}^{2}}M_{1,0}(\xi)\,d\xi$
$\leq C\|\varphi\|_{H^{\sigma}(\mathbb{R}^{d}\times\mathbb{R}^{d})}.$
Hence, we have
$|III|\leq C\|\varphi\|_{H^{\sigma}}\|u\|_{L^{\infty}}\|\rho_{f}-\rho\|_{L^{1}}+C\|\varphi\|_{H^{\sigma}}\|\rho_{f}-\rho\|_{H_{x}^{-\sigma}}.$
Combining the bounds for $I$, IIII, and $III$, and taking the supremum over all $\varphi$ with $\|\varphi\|_{H^{\sigma}}\leq 1$, we obtain
$\|f-M_{\rho,u}\|_{H_{x,\xi}^{-\sigma}}^{2}$ $\leq C\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\frac{1}{f}|\nabla_{\xi}f+(\xi-u_{f})f|^{2}\,dxd\xi+C\int_{\mathbb{R}^{d}}\rho_{f}|u_{f}-u|^{2}\,dx$
$\quad+C\|u\|_{L^{\infty}((0,T)\times\mathbb{R}^{d})}^{2}\|\rho_{f}-\rho\|_{L^{1}}^{2}+C\|\rho_{f}-\rho\|_{H_{x}^{-\sigma}}^{2}.$
Integrating over $(0,t)$ yields the desired estimate. The case $u=0$ follows immediately. ∎
@@ 3.2. Modulated energy for the Riesz interaction
A central ingredient in the relative entropy framework is the control of the nonlocal interaction term. To this end, we employ modulated interaction energies, which provide a quantitative way of comparing two densities evolving under continuity equations. Such estimates were developed in [25, Lemma 2.1] (see also [28, 53]), and we recall the version adapted to our setting below.
@@ 3.3. Relative entropy estimate
$\quad=-\frac{\rm B}{{\rm A}}\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}f\left((u-\xi)\otimes(u-\xi)-\mathbb{I}_{d}\right):\nabla u\,dxd\xi-\frac{1}{{\rm A}\tau}\int_{\mathbb{R}^{d}}\rho_{f}u_{f}\cdot(u_{f}-u)\,dx$
@@
$\quad-\frac{1}{{\rm A}\tau}\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\frac{1}{f}|\nabla_{\xi}f+(\xi-u_{f})f|^{2}dxd\xi-\frac{1}{{\rm A}\tau}\int_{\mathbb{R}^{d}}\rho_{f}u_{f}\cdot(u_{f}-u)\,dx$
@@
$II_{1}=\frac{\rm B}{\rm A}\int_{\mathbb{R}^{d}}\frac{\rho_{f}}{\rho}\nabla\cdot(\rho u)\,dx=\frac{\rm B}{\rm A}\int_{\mathbb{R}^{d}}\rho_{f}u\cdot\nabla\log\rho\,dx+\frac{\rm B}{\rm A}\int_{\mathbb{R}^{d}}\rho_{f}\mathbb{I}_{d}:\nabla u\,dx$
@@
$=-\frac{\rm B}{{\rm A}}\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}(u-\xi)\otimes uf:\nabla u\,dxd\xi-\frac{\rm B}{{\rm A}}\int_{\mathbb{R}^{d}}\rho_{f}(u-u_{f})\cdot\nabla\log\rho\,dx$
@@
$II$ $=\frac{\rm B}{\rm A}\int_{\mathbb{R}^{d}}\rho_{f}u_{f}\cdot\nabla\log\rho\,dx+\frac{\rm B}{\rm A}\int_{\mathbb{R}^{d}}\rho_{f}\mathbb{I}_{d}:\nabla u\,dx-\frac{\rm B}{{\rm A}}\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}(u-\xi)\otimes uf:\nabla u\,dxd\xi$
@@
$-\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\left(f(u-\xi)\otimes(u-\xi)-\rho_{f}\mathbb{I}_{d}\right):\nabla u\,dxd\xi$
@@ 4. Quantitative limits: Diffusive, High-Field, and gSQG
We now apply the quantitative stability tools developed in Section 3 to three hydrodynamic regimes of the kinetic model: the diffusive, high-field, and gSQG limits. Each regime corresponds to a different balance between transport, interaction, and relaxation, yet all can be analyzed within the same entropy-dissipation framework. The common strategy is to reformulate the limit equation in conservative form so that it fits into the structure of Proposition 3.2, thereby enabling a comparison with the kinetic dynamics via relative entropy. In this way, the relative entropy method is coupled with modulated interaction energies and, when appropriate, bounded Lipschitz estimates, so that microscopic dissipation can be transferred into quantitative control of macroscopic variables. Depending on the regime, this leads either to strong convergence estimates under well-prepared data or to quantitative stability in weaker topologies adapted to the corresponding limiting dynamics.
@@ 4.1. Diffusive limit
Formally, as stated in Section 1.2.1, one obtains the nonlinear diffusion equation as $\varepsilon\to 0$:
@@
$\left|\int_{\mathbb{R}^{d}}\rho^{\varepsilon}({\rm u_{\varepsilon}}-u^{\varepsilon})\cdot{\rm e}_{\varepsilon}\,dx\right|\leq\frac{1}{2\varepsilon^{2}}\int_{\mathbb{R}^{d}}\rho^{\varepsilon}|u^{\varepsilon}-{\rm u_{\varepsilon}}|^{2}\,dx+\frac{\varepsilon^{2}}{2}\|{\rm e}_{\varepsilon}\|_{L^{\infty}}\|\rho^{\varepsilon}\|_{L^{1}}\leq\frac{1}{2\varepsilon^{2}}\int_{\mathbb{R}^{d}}\rho^{\varepsilon}|u^{\varepsilon}-{\rm u_{\varepsilon}}|^{2}\,dx+C\varepsilon^{3}.$
@@
$\mathscr{H}[f^{\varepsilon}|M_{\rho,{\rm u_{\varepsilon}}}]\leq 2\mathscr{H}[f^{\varepsilon}|M_{\rho,0}]+C\varepsilon^{2}$ (4.6)
@@
$\mathscr{H}[f^{\varepsilon}|M_{\rho,0}]\leq 2\mathscr{H}[f^{\varepsilon}|M_{\rho,{\rm u_{\varepsilon}}}]+C\varepsilon^{2}.$
@@
Estimate (4.7), together with Lemma 2.2, ensures the convergence of $f^{\varepsilon}$ and $\rho^{\varepsilon}$. For the momentum, we note
@@
We next derive the sharper low-topology estimate for $\rho^{\varepsilon}$, which is consistent with the optimal diffusive scaling suggested by the formal expansion in Appendix B.1. To this end, we assume that either $T$ is sufficiently small or $\|\rho\|_{L^{\infty}(0,T;H^{\beta+1})}$ is sufficiently small for some $\beta>\frac{d}{2}+1$.
Recall that, by Definition 1.1, $f^{\varepsilon}$ satisfies the kinetic equation in the sense of distributions. In particular, the force term $\nabla(-\Delta)^{-\alpha}\rho^{\varepsilon}\cdot\nabla_{\xi}f^{\varepsilon}$ is interpreted through integration by parts in the velocity variable, which implicitly requires
$f^{\varepsilon}\,\nabla(-\Delta)^{-\alpha}\rho^{\varepsilon}\in L^{1}_{\mathrm{loc}}((0,T)\times\mathbb{R}^{d}\times\mathbb{R}^{d}).$
Since the force field does not depend on $\xi$, this implies
$\rho^{\varepsilon}\,\nabla(-\Delta)^{-\alpha}\rho^{\varepsilon}\in L^{1}_{\mathrm{loc}}((0,T)\times\mathbb{R}^{d}),$
so that the macroscopic drift term
$\nabla\cdot(\rho^{\varepsilon}\nabla(-\Delta)^{-\alpha}\rho^{\varepsilon})$
is well defined in the sense of distributions.
Moreover, since
$f^{\varepsilon}\in L^{\infty}\bigl(0,T;L^{1}(\mathbb{R}^{d}\times\mathbb{R}^{d},(1+|\xi|^{2})\,dxd\xi)\bigr),$
the velocity moments
$\rho^{\varepsilon}:=\int_{\mathbb{R}^{d}}f^{\varepsilon}\,d\xi,\quad m^{\varepsilon}:=\int_{\mathbb{R}^{d}}\xi f^{\varepsilon}\,d\xi=\rho^{\varepsilon}u^{\varepsilon},\quad S^{\varepsilon}:=\int_{\mathbb{R}^{d}}(\xi\otimes\xi-\mathbb{I}_{d})f^{\varepsilon}\,d\xi$
belong to $L^{\infty}(0,T;L^{1}(\mathbb{R}^{d}))$, $L^{\infty}(0,T;L^{1}(\mathbb{R}^{d};\mathbb{R}^{d}))$, and $L^{\infty}(0,T;L^{1}(\mathbb{R}^{d};\mathbb{R}^{d\times d}))$, respectively. Hence, by integrating the kinetic equation with respect to $\xi$, we deduce that $\rho^{\varepsilon}$ satisfies
$\partial_{t}\rho^{\varepsilon}-\Delta\rho^{\varepsilon}=\nabla\cdot(\rho^{\varepsilon}\nabla(-\Delta)^{-\alpha}\rho^{\varepsilon})+\varepsilon\nabla\cdot\partial_{t}m^{\varepsilon}+\nabla\otimes\nabla:S^{\varepsilon}$
in the sense of distributions on $(0,T)\times\mathbb{R}^{d}$.
We work with the inhomogeneous Sobolev space $H^{\beta}(\mathbb{R}^{d})$ equipped with the norm
$\|f\|_{H^{\beta}}^{2}:=\int_{\mathbb{R}^{d}}(1+|\eta|^{2})^{\beta}|\hat{f}(\eta)|^{2}\,d\eta.$
Its dual space is $H^{-\beta}(\mathbb{R}^{d})$. Since $L^{1}(\mathbb{R}^{d})\subset H^{-s}(\mathbb{R}^{d})$ for any $s>\frac{d}{2}$, and $\beta>\frac{d}{2}+1$, we have
$\rho^{\varepsilon}(t),\,\rho(t)\in L^{1}(\mathbb{R}^{d})\subset H^{-\beta}(\mathbb{R}^{d})\quad\text{for all }t\in[0,T].$
Therefore the duality pairing $\langle\rho^{\varepsilon}(t)-\rho(t),\varphi\rangle$ is well defined for any $\varphi\in H^{\beta}(\mathbb{R}^{d})$.
We now record the corresponding Duhamel formula in duality form. For any $\varphi\in H^{\beta}(\mathbb{R}^{d})$, one has
$\langle\rho^{\varepsilon}(t),\varphi\rangle$ $=\langle\rho_{0}^{\varepsilon},\mathcal{H}_{t}\star_{x}\varphi\rangle-\int_{0}^{t}\Big\langle\rho^{\varepsilon}\nabla(-\Delta)^{-\alpha}\rho^{\varepsilon},\,\nabla(\mathcal{H}_{t-s}\star_{x}\varphi)\Big\rangle\,ds$
$\quad-\varepsilon\langle\rho^{\varepsilon}u^{\varepsilon}(t),\nabla\varphi\rangle+\varepsilon\langle\rho_{0}^{\varepsilon}u_{0}^{\varepsilon},\nabla(\mathcal{H}_{t}\star_{x}\varphi)\rangle$
$\quad-\varepsilon\int_{0}^{t}\langle\rho^{\varepsilon}u^{\varepsilon}(s),\nabla\Delta(\mathcal{H}_{t-s}\star_{x}\varphi)\rangle\,ds+\int_{0}^{t}\langle S^{\varepsilon}(s),\nabla^{2}(\mathcal{H}_{t-s}\star_{x}\varphi)\rangle\,ds,$
where $\mathcal{H}_{t}(x):=(4\pi t)^{-d/2}e^{-\frac{|x|^{2}}{4t}}$ is the heat kernel and $\star_{x}$ denotes the convolution with respect to $x$. Here we used
$\int_{0}^{t}f(t-s)\frac{dg}{ds}(s)\,ds=\int_{0}^{t}\frac{df}{ds}(t-s)g(s)\,ds+f(0)g(t)-f(t)g(0)$
and $\partial_{s}(\mathcal{H}_{s}\star_{x}f)=\Delta(\mathcal{H}_{s}\star_{x}f)$. The same identity holds for $\rho$, with the $m^{\varepsilon}$- and $S^{\varepsilon}$-terms removed.
Now, let $\varphi\in H^{\beta}(\mathbb{R}^{d})$ be arbitrary. Applying the above duality formula to $\rho^{\varepsilon}$ and $\rho$, and taking the difference, we obtain
$\int_{\mathbb{R}^{d}}(\rho^{\varepsilon}-\rho)\varphi\,dx$
$\quad=\int_{\mathbb{R}^{d}}\mathcal{H}_{t}\star_{x}(\rho_{0}^{\varepsilon}-\rho_{0})\varphi\,dx-\int_{0}^{t}\int_{\mathbb{R}^{d}}(\rho^{\varepsilon}-\rho)(s)\nabla(-\Delta)^{-\alpha}(\rho^{\varepsilon}-\rho)(s)\mathcal{H}_{t-s}\star_{x}\nabla\varphi\,dxds$
$\qquad-\int_{0}^{t}\int_{\mathbb{R}^{d}}(\rho^{\varepsilon}-\rho)(s)\left[(\nabla(-\Delta)^{-\alpha}\rho)(s)\cdot(\mathcal{H}_{t-s}\star_{x}\nabla\varphi)-\nabla\cdot(-\Delta)^{-\alpha}(\rho(s)\mathcal{H}_{t-s}\star_{x}\nabla\varphi)\right]\,dxds$